# Strong Screening Rules for Group-based SLOPE Models

### v1

@@ 3 Group SLOPE
The strong screening rule for gSLOPE presented in this section is formulated by computing the zero condition of its subdifferential (as per Equation 3) derived in Theorem 3.1. To derive the subdifferential, we define the operator
@@
$\partial J_{\text{gslope}}(\beta;w)=\begin{dcases}\{x\in\mathbb{R}^{\operatorname{card}{\mathcal{G}_{\mathcal{Z}}}}:{[x]}_{\mathcal{G}_{\mathcal{Z}},-0.5}\in\partial J_{\text{slope}}(0;w_{\mathcal{Z}})\},&\text{at 0.}\\ w_{g}\sqrt{p_{g}}\frac{\beta^{(g)}}{\|\beta^{(g)}\|_{2}},&\text{otherwise.}\end{dcases}$
The choice of $q=-0.5$ leads to $J_{\text{gslope}}^{*}(x;w)=J_{\text{slope}}^{*}([x]_{\mathcal{G},-0.5})$, which allows the gSLOPE subdifferential to be written in terms of the SLOPE one [6]. Combining the KKT conditions (Equation 2) with the gSLOPE subdifferential (Theorem 3.1) reveals that a group is zero if $h(\lambda):=([\nabla f(\hat{\beta}(\lambda))]_{\mathcal{G},-0.5})_{\downarrow}\in\partial J_{\text{slope}}(\mathbf{0};\lambda w_{\mathcal{Z}})$. Using the subdifferential of SLOPE [17], this is given by
$\text{cumsum}(h(\lambda)-\lambda w_{\mathcal{Z}})\preceq\mathbf{0}.$ (5)
This condition is checked efficiently using the algorithm proposed for the SLOPE strong rule (Algorithm A1) leading to the strong rule for gSLOPE (Proposition 3.2). The algorithm assumes that the indices for the inactive predictors will be ordered last in the input $c$ and the features $|\hat{\beta}|_{\downarrow}$ [17].
@@
The gradient at path value $k+1$ is not available for the computation of $h(\lambda_{k+1})$, so an approximation is required that does not lead to any violations in Algorithm A1. By the cumsum condition in this algorithm, an approximation for a group $g$ is sought such that $h_{g}(\lambda_{k+1})\leq h_{g}(\lambda_{k})+R_{g}$, where $R_{g}\geq 0$ needs to be determined. An approximation is found by assuming that $h_{g}(\lambda_{k+1})$ is a Lipschitz function of $\lambda_{k+1}$ with respect to the $\ell_{1}$ norm, that is,
@@
By again noting that $J_{\text{gslope}}^{*}(x)=J_{\text{slope}}^{*}([x]_{\mathcal{G},-0.5})$, it can be seen that the assumption is equivalent to those used to derive strong rules for the lasso and SLOPE [17, 31]. By the reverse triangle inequality,
@@
leading to the choice $R_{g}=\lambda_{k}w_{g}-\lambda_{k+1}w_{g}$ and the gradient approximation strong screening rule.
@@
Taking $c=h(\lambda_{k})+\lambda_{k}w-\lambda_{k+1}w$ and $\phi=\lambda_{k+1}w$ as inputs for Algorithm A1, and assuming that for any $k\in\{1,\dots,l-1\}$,
$\left|h_{g}(\lambda_{k+1})-h_{g}(\lambda_{k})\right|\leq w_{g}|\lambda_{k+1}-\lambda_{k}|,\;\forall g=1,\dots,m,$
@@ 4 Sparse-group SLOPE
$-\nabla f(\beta)\in\lambda\alpha\partial J_{\text{slope}}(\beta;v)+\lambda(1-\alpha)\partial J_{\text{gslope}}(\beta;w).$ (6)
@@ 4.1 Group screening
The problem reduces to that of the gSLOPE screening rule (Section 3), with input given by $c=([\nabla f(\beta)+\lambda\alpha\partial J_{\text{slope}}(\mathbf{0};v)]_{\mathcal{G},-0.5})_{\downarrow}$ and $\phi=\lambda(1-\alpha)w$. To determine the form of the quantity $\partial J_{\text{slope}}(\mathbf{0};v)$, the term inside the $[\cdot]_{\mathcal{G}_{\mathcal{Z}},-0.5}$ operator needs to be as small as possible for Equation 7 to be satisfied. This term is found to be the soft thresholding operator, $S(\nabla f(\beta),\lambda\alpha):=\text{sign}(\nabla f(\beta))(|\nabla f(\beta)|-\lambda\alpha)_{+}$ (with the derivation presented in Appendix B.2).
By using the soft-thresholding operator, a valuable connection between SGS and SGL is made, as this operator is used in the gradient update step for SGL [29]. Such a connection has the potential to lead to new and more efficient optimization approaches for SGS that are more closely related to those used to solve SGL, similar to the recently developed coordinate descent algorithm for SLOPE [18].
Using this function, the (non-approximated) strong group screening rule for SGS is given in Proposition B.1. Using a similar Lipschitz assumption as for the gSLOPE rule gives the gradient approximation strong group screening rule for SGS (Proposition 4.1).
@@
Let $\tilde{h}(\lambda):=([S(\nabla f(\hat{\beta}(\lambda)),\lambda\alpha)]_{\mathcal{G},-0.5})_{\downarrow}$. Taking $c=\tilde{h}(\lambda_{k})+\lambda_{k}(1-\alpha)w-\lambda_{k+1}(1-\alpha)w$ and $\phi=\lambda_{k+1}(1-\alpha)w$ as inputs for Algorithm A1, and assuming that for any $k\in\{1,\dots,l-1\}$,
$\left|\tilde{h}_{g}(\lambda_{k+1})-\tilde{h}_{g}(\lambda_{k})\right|\leq(1-\alpha)w_{g}|\lambda_{k+1}-\lambda_{k}|,\forall g=1,\dots,m,$
and $\mathcal{O}(\tilde{h}(\lambda_{k+1}))=\mathcal{O}(\tilde{h}(\lambda_{k}))$, then the algorithm returns a superset $\mathcal{S}_{g}(\lambda_{k+1})$ of $\mathcal{A}_{g}(\lambda_{k+1})$.
@@ 4.2 Variable screening
Whilst the group screening for SGS returns a superset, $\mathcal{S}_{g}$, of the active group set, exploiting the sparse-group penalization of SGS allows for the screened set to be further reduced by also applying variable screening. The KKT conditions (Equation 6) for a zero variable, $j\in\mathcal{G}_{g}$, in an active group, $g\in\mathcal{A}_{g}$, are
$-\nabla_{j}f(\beta)\in\lambda\alpha\partial J_{\text{slope}}(\mathbf{0};v_{j}).$ (8)
The gSLOPE subdifferential term vanishes as the numerator is zero in Theorem 3.1. The problem reduces to that of SLOPE variable screening, applied only to the variables that belong to groups in $\mathcal{A}_{g}$ and scaled by $\alpha$. The gradient approximated rule is formalized in Proposition 4.2 (the non-approximated version is presented in Proposition B.2).
@@
Let $\bar{h}(\lambda)=(\nabla f(\hat{\beta}(\lambda)))_{|\downarrow|}$. Taking $c=|\bar{h}(\lambda_{k+1})|+\lambda_{k}\alpha v-\lambda_{k+1}\alpha v$ and $\phi=\lambda_{k+1}\alpha v$ for only the variables in the groups in $\mathcal{A}_{g}(\lambda_{k+1})$ as inputs for Algorithm A1 , and assuming that for any $k\in\{1,\dots,l-1\}$,
@@
In practice $\mathcal{A}_{g}(\lambda_{k+1})$ is not available, as this is exactly what we are trying to superset with any screening rule. However, Proposition 4.1 states that it is contained within $\mathcal{S}_{g}(\lambda_{k+1})$, so that the screened set can be used instead and complemented with KKT checks to ensure there are no violations, which would be performed anyway. For the checks, the group KKT conditions (Equation 6) are checked. For any active groups or groups violating the check (indicating that they should be active), the variable KKT conditions (Equation 8) are checked on the variables within those groups, to identify those that need to be added to $\mathcal{E}_{v}$.
@@ 5 Results
In this section, the effectiveness of the screening rules for gSLOPE and SGS is illustrated through the analysis of both synthetic (Section 5.1) and real data (Section 5.2). Any mentions of $\mathcal{E},\mathcal{A}$ in reference to group and variable metrics refers to $\mathcal{E}_{g},\mathcal{A}_{g}$ and $\mathcal{E}_{v},\mathcal{A}_{v}$ respectively. For SGS, $\mathcal{E}_{g}$ is the groups with representation in $\mathcal{E}_{v}$ (the groups fitted). Computational information is presented in Appendix F.1.
@@ 5.1 Synthetic data analysis
For the synthetic data, a multivariate Gaussian design matrix, $\mathbf{X}\sim\mathcal{N}(\mathbf{0},\boldsymbol{\Sigma})\in\mathbb{R}^{400\times p}$, was used and the within-group correlation set to $\Sigma_{i,j}=\rho$, where $i$ and $j$ belong to the same group. The correlation and number of features were varied between $\rho\in\{0,0.3,0.6,0.9\}$ and $p\in\{500,1625,2750,3875,5000\}$, producing $20$ simulation cases. Each simulation case was repeated $100$ times. Two models were considered: linear and logistic. For the linear model, the output was generated as $y=\mathbf{X}\beta+\mathcal{N}(0,1)$ and for the logistic model the class probabilities were calculated using $\sigma(\mathbf{X}\beta+\mathcal{N}(0,1))$, where $\sigma$ is the sigmoid function. Groups of sizes between $3$ and $25$ were considered, of which $15\%$ were set to active. Within each group, $30\%$ of the variables were set to active with the signal sampled from $\beta\sim\mathcal{N}(0,5)$.
gSLOPE and SGS were each fit along a log-linear path of $l=50$ regularization parameters using warm starts. The first path value $\lambda_{1}$ was set to the exact value at which the first group enters the model (given in Appendix A.3 for gSLOPE and Appendix B.4 for SGS). The final regularization value was set to $\lambda_{50}=0.05\lambda_{1}$. The data was $\ell_{2}$ standardized and for the linear model an intercept was fitted. Both models had FDR-control parameters set to $0.05$, and $\alpha=0.95$ for SGS. Even though the models were fit using the adaptive three operator splitting algorithm (ATOS) [26], it should be noted that the screening rules are independent of the choice of fitting algorithm.
Primarily, the results for the linear model are presented, and the corresponding plots for the logistic model are presented in Appendix F.3.2.
Screening efficiency.
By comparing the sizes of the fitting set $(\mathcal{E})$ to the active set ($\mathcal{A})$, we can understand the difference between the minimum dimensionality needed against the dimensionality provided by the screening rules. As expected, these sizes increase as $\lambda$ decreases (Figure 2). In particular, the gap between sets widens as more features enter the model. The size of $\mathcal{E}$ remains far below the size of the full input space even at the lowest value of $\lambda$, showing that the screening rules have a benefit along the whole path. This finding is found to be independent of $\rho$, $p$, and model fitted (see Appendix F.3).
@@
Figure 2: The number of groups/variables in $\mathcal{E},\mathcal{A}$ for both gSLOPE and SGS as a function of the regularization path for the linear model with $p=2750,\rho=0.6,m=197$. The results are averaged over $100$ repetitions, with the shaded regions corresponding to $95\%$ confidence intervals.
The performance of the screening rules are generally similar across linear and logistic models (Figure 3). As the within-group correlation increases, the size of both the screened and active sets decreases, causing the variables to move in the same direction within the same group. Due to the shape of the convex penalty of SLOPE, highly correlated features are clustered together [36]. This leads to smaller active sets, and in turn, also smaller screened sets. Additionally, it can be seen that the size difference between the screened and active sets decreases with correlation. Across the different correlation values, the linear model had smaller set sizes than the logistic one.
Runtime performance.
A key metric of performance for a screening rule is the time taken to fit a path of models. Increasing the value of $p$ demonstrates the computational cost reduction of applying a screening rule (Figure 4). For the smaller values of $p$, the runtime tends to be similar whether screening is applied or not. However, once the input dimension increases, the benefit of screening can be seen clearly. Additionally, increasing the within-group correlation tends to lead to smaller active and screened sets, but the benefit of screening appears to be robust under different correlations.
Figure 5: Runtime (in seconds) for fitting $100$ models along a path for screening against no screening for gSLOPE and SGS, averaged across the nine collections of pathways for each dataset, with standard errors shown.
A clear improvement with regards to runtime can be seen aggregating the results across all simulation cases (Table 1). In particular, for both models considered, screening halves the runtime for fitting $50$ models for SGS.
5.2 Real data experiments
@@
Table 2: The sizes of the sets $\mathcal{A}$ and $\mathcal{E}$ (with standard errors), averaged across $100$ path values and pathway collections for each of the two datasets. SGS var/grp shows the corresponding metrics for $\mathcal{A}_{v},\mathcal{E}_{v}$ and $\mathcal{A}_{g},\mathcal{E}_{g}$ for the bi-level screening procedure.
Method Dataset $\operatorname{card}(\mathcal{A})$ $\operatorname{card}(\mathcal{E})$
gSLOPE Cancer $66\pm 1$ $120\pm 2$
gSLOPE Colitis $32\pm 1$ $56\pm 2$
SGS grp Cancer $50\pm 1$ $115\pm 2$
SGS grp Colitis $18\pm 1$ $87\pm 3$
SGS var Cancer $71\pm 1$ $212\pm 4$
SGS var Colitis $25\pm 1$ $209\pm 6$
Datasets.
The screening rules were applied to two gene expression datasets with binary responses. The datasets were previously analyzed to measure the predictive performance of sparse-group regression models on large datasets [14, 29]. The cancer dataset contains gene data for $60$ breast cancer patients who had been treated with tamoxifen for 5 years, classified into binary labels depending on whether the cancer recurred [20]. The colitis dataset uses transcriptional profiles in peripheral blood mononuclear cells of $127$ patients to determine which patients have an inflammatory bowel disease (ulcerative colitis or Crohn’s). Of the $127$ samples, $42$ of them are controls [7].
The genes of both datasets were assigned to pathways (groups) that were obtained from the Molecular Signature Database. [^1] https://www.gsea-msigdb.org/gsea/msigdb/human/collections.jsp Nine collections of pathways were downloaded and as the pathways included different genes, nine unique design matrices for each dataset were created, each with a unique grouping structure. This gave $18$ design matrices in total to which the screening rules were applied.

### v2

@@ 3 GROUP SLOPE
The strong rule for gSLOPE is formulated by checking the zero condition of the subdifferential (as per Equation 3) derived in Theorem 3.1. To derive the subdifferential, we define the operator
@@
$\partial J_{\text{gslope}}(\beta;w)=\begin{dcases}\begin{array}[]{l}\big{\{}x\in\mathbb{R}^{\operatorname{card}\mathcal{G}_{\mathcal{Z}}}:\\ {[x]}_{\mathcal{G}_{\mathcal{Z}},-0.5}\in\partial J_{\text{slope}}(0;w_{\mathcal{Z}})\big{\}},\end{array}\text{at 0.}&\\ \left\{w_{g}\sqrt{p_{g}}\frac{\beta^{(g)}}{\|\beta^{(g)}\|_{2}}\right\},\;\;\;\;\;\;\;\;\;\text{otherwise.}&\end{dcases}$
The choice of $q=-0.5$ leads to $J_{\text{gslope}}^{*}(x;w)=J_{\text{slope}}^{*}([x]_{\mathcal{G},-0.5})$, which allows the gSLOPE subdifferential to be written in terms of the SLOPE one (Brzyski et al., 2019). Combining the KKT conditions at zero (Equation 3) with the gSLOPE subdifferential (Theorem 3.1) reveals that a group is inactive if $h(\lambda):=([\nabla f(\hat{\beta}(\lambda))]_{\mathcal{G},-0.5})_{\downarrow}\in\partial J_{\text{slope}}(\mathbf{0};\lambda w)$. Using the subdifferential of SLOPE (Appendix A.2) (Larsson et al., 2020), this is given by
$\text{cumsum}(h(\lambda)-\lambda w)\preceq\mathbf{0}.$ (5)
This condition is checked efficiently using the algorithm proposed for the SLOPE strong rule (Algorithm A1) leading to the strong rule for gSLOPE (Proposition 3.2). The algorithm assumes that the indices for the inactive predictors will be ordered last in the input $c$ and the features $|\hat{\beta}|_{\downarrow}$ (Larsson et al., 2020).
@@
The gradient at path index $k+1$ is not available for the computation of $h(\lambda_{k+1})$, so an approximation is required that does not lead to any violations in Algorithm A1. By the cumsum condition in this algorithm, an approximation for a group $g\in[m]$ is sought such that $h_{g}(\lambda_{k+1})\leq h_{g}(\lambda_{k})+R_{g}$, where $R_{g}\geq 0$ needs to be determined. An approximation is found by assuming that $h_{g}(\lambda_{k+1})$ is a Lipschitz function of $\lambda_{k+1}$ with respect to the $\ell_{1}$ norm, that is,
@@
By again noting that $J_{\text{gslope}}^{*}(x)=J_{\text{slope}}^{*}([x]_{\mathcal{G},-0.5})$, it can be seen that the assumption is equivalent to the Lipschitz assumptions used for the lasso and SLOPE strong rules (Tibshirani et al., 2010; Larsson et al., 2020). By the reverse triangle inequality,
@@
leading to the choice $R_{g}=\lambda_{k}w_{g}-\lambda_{k+1}w_{g}$ and the gradient approximation strong screening rule (Proposition 3.3). To apply gSLOPE screening in practice (Section 6.1), Proposition 3.4 describes the calculation of the first path value.
@@
Taking $c=h(\lambda_{k})+\lambda_{k}w-\lambda_{k+1}w$ and $\phi=\lambda_{k+1}w$ as inputs for Algorithm A1, and assuming that for any $k\in[l-1]$,
$\left|h_{g}(\lambda_{k+1})-h_{g}(\lambda_{k})\right|\leq w_{g}|\lambda_{k+1}-\lambda_{k}|,\;\forall g=[m],$
@@
Proposition 3.4 (gSLOPE path start).
For gSLOPE, the path value at which the first group enters the model is given by
$\lambda=\max\left\{\operatorname{cumsum}\bigl{(}([\nabla f(\mathbf{0})]_{\mathcal{G},-0.5})_{\downarrow}\bigr{)}\oslash\operatorname{cumsum}(w)\right\},$
[…]
@@ 4 SPARSE-GROUP SLOPE
$\nabla f(\beta)\in\lambda\alpha\partial J_{\text{slope}}(\beta;v)+\lambda(1-\alpha)\partial J_{\text{gslope}}(\beta;w).$ (6)
@@ 4.1 Group Screening
$\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;-\lambda(1-\alpha)w_{\mathcal{Z}}\Bigr{)}\preceq\mathbf{0}.$ (7)
The problem reduces to a form similar to the gSLOPE screening rule (Section 3), with inputs for Algorithm A1 given by $c=([\nabla f(\beta)+\lambda\alpha\partial J_{\text{slope}}(\mathbf{0};v)]_{\mathcal{G},-0.5})_{\downarrow}$ and $\phi=\lambda(1-\alpha)w$.
To determine the form of the quantity $\partial J_{\text{slope}}(\mathbf{0};v)$, the term inside the $[\cdot]_{\mathcal{G}_{\mathcal{Z}},-0.5}$ operator needs to be as small as possible for Equation 7 to be satisfied. This term is found to be the soft thresholding operator, $S(\nabla f(\beta),\lambda\alpha v):=\text{sign}(\nabla f(\beta))(|\nabla f(\beta)|-\lambda\alpha v)_{+}$ by Lemma 4.1 (see Appendix B.2 for the proof).
Lemma 4.1.
In Equation 7, choosing $\partial J_{\text{slope}}(\mathbf{0};v)=S(\nabla f(\beta),\lambda\alpha v)$ minimises $[\nabla f(\beta)+\lambda\alpha\partial J_{\text{slope}}(\mathbf{0};v)]_{\mathcal{G}}$.
By using the soft-thresholding operator, a valuable connection between SGS and SGL is found, as the operator is used in the gradient update step for SGL (Simon et al., 2013). This connection has the potential to lead to new and more efficient optimization approaches for SGS that are more closely related to those used to solve SGL, similar to the recent coordinate descent algorithm for SLOPE (Larsson et al., 2022).
Using this operator, the (non-approximated) strong group screening rule for SGS is shown in Proposition B.1. Applying a similar Lipschitz assumption as for the gSLOPE rule gives the gradient approximation strong group screening rule for SGS (Proposition 4.2).
@@
Let $\tilde{h}(\lambda):=([S(\nabla f(\hat{\beta}(\lambda)),\lambda\alpha v)]_{\mathcal{G},-0.5})_{\downarrow}$. Taking $c=\tilde{h}(\lambda_{k})+\lambda_{k}(1-\alpha)w-\lambda_{k+1}(1-\alpha)w$ and $\phi=\lambda_{k+1}(1-\alpha)w$ as inputs for Algorithm A1, and assuming that for any $k\in[l-1]$,
$\left|\tilde{h}_{g}(\lambda_{k+1})-\tilde{h}_{g}(\lambda_{k})\right|\leq(1-\alpha)w_{g}|\lambda_{k+1}-\lambda_{k}|,\;\forall g=[m],$
and $\mathcal{O}(\tilde{h}(\lambda_{k+1}))=\mathcal{O}(\tilde{h}(\lambda_{k}))$, then the algorithm returns a superset $\mathcal{S}_{g}(\lambda_{k+1})$ of the active set $\mathcal{A}_{g}(\lambda_{k+1})$.
@@ 4.2 Variable Screening
By exploiting the sparse-group norm of SGS, the input dimensionality can be reduced further with a second layer of variable screening. The KKT conditions (Equation 6) for zero variables in active groups are
$\nabla_{\mathcal{G}_{\mathcal{A}_{g}}}f(\beta)\in\lambda\alpha\partial J_{\text{slope}}(\mathbf{0};v_{\mathcal{G}_{\mathcal{A}_{g}}}).$ (8)
The gSLOPE subdifferential term vanishes as the numerator is zero in Theorem 3.1. The problem reduces to that of SLOPE screening, applied to the variables in groups in $\mathcal{A}_{g}$ and scaled by $\alpha$. The gradient approximated rule is shown in Proposition 4.3 (see Proposition B.2 for the non-approximated version).
@@
Let $\bar{h}(\lambda):=(\nabla f(\hat{\beta}(\lambda)))_{|\downarrow|}$. Taking $c=|\bar{h}(\lambda_{k+1})|+\lambda_{k}\alpha v-\lambda_{k+1}\alpha v$ and $\phi=\lambda_{k+1}\alpha v$ for the variables in the groups in $\mathcal{A}_{g}(\lambda_{k+1})$ as inputs for Algorithm A1, and assuming that for any $k\in[l-1]$,
@@
In practice $\mathcal{A}_{g}(\lambda_{k+1})$ is not available, as this is exactly what we are trying to superset with any screening rule. However, Proposition 4.2 guarantees that it is contained in $\mathcal{S}_{g}(\lambda_{k+1})$ so that this can be used instead. To apply SGS in practice (Section 6.1), Proposition 4.4 describes the calculation of the first path value.
Proposition 4.4 (SGS path start).
For SGS, the path value at which the first variable enters the model is
$\lambda=\max\{\operatorname{cumsum}(|\nabla$ $f(\mathbf{0})|_{\downarrow})\oslash$
$\operatorname{cumsum}((1-\alpha)\tau\omega-\alpha v)\},$
where $\tau$ and $\omega$ are expanded vectors of the group sizes ($\sqrt{p_{g}})$ and penalty weights ($w_{g}$) to $p$ dimensions.
5 GROUP-BASED OWL
The screening rule framework presented are also applicable to the wider class of OWL models. The Ordered Weighted $\ell_{1}$ (OWL) framework is defined as (Zeng and Figueiredo, 2014a)
$\hat{\beta}\in\operatorname*{arg\,min}_{\beta\in\mathbb{R}^{p}}\left\{\nabla f(\beta)+\lambda J_{\text{owl}}(\beta;v)\right\},$
where $J_{\text{owl}}(\beta;v)=\sum_{i=1}^{p}v_{i}|\beta|_{(i)}$, $|\beta|_{(1)}\geq\ldots\geq|\beta|_{(p)}$, and $v$ are non-negative non-increasing weights. SLOPE is a special case of OWL where the weights are taken to be the Benjamini-Hochberg critical values (Bogdan et al., 2015). Octagonal Shrinkage and Clustering Algorithm for Regression (OSCAR) (Bondell and Reich, 2008) is a further special case of OWL (often referred to as OWL with linear decay) where for a variable $i\in[p]$, the weights are taken to be $v_{i}=\sigma_{1}+\sigma_{2}(p-i)$, and $\sigma_{1},\sigma_{2}$ are to be set. In Bao et al. (2020) they are set to $\sigma_{1}=d_{i}\|\mathbf{X}^{\top}y\|_{\infty},\sigma_{2}=\sigma_{1}/p$, where $d_{i}=i\times e^{-2}$, $\mathbf{X}^{\top}\in\mathbb{R}^{n\times p}$ is the design matrix, and $y\in\mathbb{R}^{n}$ is the response vector.
Group OSCAR (gOSCAR) and Sparse-group OSCAR (SGO) are defined using the frameworks provided by gSLOPE (Brzyski et al., 2019) and SGS (Feser and Evangelou, 2023), respectively, but instead use the OSCAR weights (see Appendix E.1).
@@ 6 RESULTS
This section illustrates the effectiveness of the screening rules for gSLOPE and SGS using synthetic (Section 6.1) and real (Section 6.2) data. References to $\mathcal{E},\mathcal{A}$ for group and variable metrics denote $\mathcal{E}_{g},\mathcal{A}_{g}$ and $\mathcal{E}_{v},\mathcal{A}_{v}$. For SGS, $\mathcal{E}_{g}$ represents groups with members in $\mathcal{E}_{v}$.
@@ 6.1 Synthetic Data Analysis
A multivariate Gaussian design matrix, $\mathbf{X}\sim\mathcal{N}(\mathbf{0},\boldsymbol{\Sigma})\in\mathbb{R}^{400\times p}$, was used and the within-group correlation set to $\Sigma_{i,j}=\rho$, where $i$ and $j$ belong to the same group. The correlation and number of features were varied between $\rho\in\{0,0.3,0.6,0.9\}$ and $p\in\{500,1625,2750,3875,5000\}$, producing $20$ simulation cases. Each simulation case was repeated $100$ times. Two models were considered: linear and logistic. For the linear model, the output was generated as $y=\mathbf{X}\beta+\mathcal{N}(0,1)$ and for the logistic model the class probabilities were calculated using $\sigma(\mathbf{X}\beta+\mathcal{N}(0,1))$, where $\sigma$ is the sigmoid function. Groups of sizes between $3$ and $25$ were considered, of which $15\%$ were set to active. Within each active group, $30\%$ of the variables were set to active with signal $\beta\sim\mathcal{N}(0,5)$.
@@
Figure 2: The proportion of groups/variables in $\mathcal{E},\mathcal{A}$, relative to the input, for both gSLOPE and SGS as a function of the path for the linear model with $p=2750,\rho=0.6,m=197$. The results are averaged over $100$ repetitions, with $95\%$ confidence intervals shown.
gSLOPE and SGS were fit along a log-linear path of $50$ regularization parameters using warm starts, beginning at $\lambda_{1}$ (from Propositions 3.4 and 4.4), and ending at $\lambda_{50}=0.05\lambda_{1}$. The data was $\ell_{2}$ standardized and for the linear model an intercept was used. Both models had FDR-control parameters set to $0.05$, and $\alpha=0.95$ for SGS. The models were fit using the adaptive three operator splitting (ATOS) algorithm (Pedregosa and Gidel, 2018), although the screening rules can be applied with any fitting algorithm. Additional computational details are in Appendix F.1.
Primarily, the results for the linear model are presented, and the results for the logistic model are in Appendix F.3.2. The simulations were repeated for group-based OSCAR models (Appendix E).
@@
Screening Efficiency
By comparing the sizes of the fitting set $(\mathcal{E})$ to the active set ($\mathcal{A})$, the screening rules are found to be efficient in providing dimensionality reduction close to the minimum possible (the active set size) (Figures 2 and A9). As expected, the sets increase in size as $\lambda$ decreases, and the difference in size between the sets remains stable along the path, decreasing towards the termination point. The size of the fitting set remains far below the input size across the whole path, showing the benefit of the screening. This is found to be true for any correlation, input dimensionality, and model considered (Appendix F.3).
The screening rules perform well for linear and logistic models (Figure 3), showing robust dimensionality reduction for all correlation cases considered. As the correlation increases, the signal concentrates in fewer groups, causing the active group set to decrease in size. SLOPE models deal well with highly correlated features, as the sorted norm clusters them together (Zeng and Figueiredo, 2014b).
The screening rules are found to efficiently reduce the input dimensionality on average across all cases considered (Table 2).
Runtime Performance
A key metric of performance for a screening rule is the time taken to fit a path of models. Figure 4 shows the significant runtime improvements our screening rules provide as a function of increasing input dimensionality. The gain is observed to be substantial under all correlation cases. Applying screening improves the scaleability of applying gSLOPE and SGS models to larger datasets.
The clear benefit and robustness of our screening method, with regards to runtime, can be seen by aggregating the results of all the simulation cases (Table 1). For both models, in the linear and logistic cases, screening substantially improves the computational cost, halving the runtime for SGS models.
6.2 Real Data Experiments
Datasets
The screening rules were applied to seven real gene datasets, of different response types and dimensionality. Two of the datasets, carbotax and sheetz, had a continuous response so were fit using a linear model. For these, the groups were generated using K-means clustering (Lloyd, 1982). The remaining five (adenoma, cancer, celiac, colitis, and tumour), had binary labels, so a logistic model was used. For these, the design matrices contained gene expression data downloaded from NCBI’s GEO database (Edgar et al., 2002), so the genes could be assigned to pathways (groups) using the C3 regulatory target gene sets from MSigDB (Subramanian et al., 2005; Liberzon et al., 2011). All datasets were high-dimensional. See Appendix F.4 for full details.